OpenStudy (anonymous):

determine the greatest possible number of elements of a subset A of {1,2,.....,30} with the property that any product of one or more distinct elements of A is not a square.

OpenStudy (anonymous):

how many elements,i think the question is asking

OpenStudy (anonymous):

how any numbers in the set , are not squares when multiplied

OpenStudy (anonymous):

actually this question do they also mean we can also multiply the number by its self or we muldiply different numbers to check if their product is a squre?

OpenStudy (anonymous):

can we say all primes are never perfect squares when multiplied by a natural number, ie <30 \[p \times n=c^2\] eg 3(n)=perfect square

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

oh yah but it has to be multiplied by its cube,so how d we eliminate non squares